ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: 41A30, 65D10, 41A25, 33E30 Volk, Wolfgang, Berlin Quasi-interpolation with Arbitrarily Often Differentiable Functions The ideas of quasi-interpolation, which often is also referred to by local approximation method, have grown up first in the context of spline functions. This approach may also be applied to other finite sets of real-valued functions satisfying certain conditions. The lecture will be concerned with quasiinterpolation with sets of arbitrarily often differentiable functions, which have been introduced by the author (cf. : Properties of subspaces generated by an infinitely often differentiable function and its translates, ZAMM Z. angew. Math. Mech. 76 (1996) Suppl. 1, 575 - 576). Primarily error estimates will be developed, but results on numerical experiments will be presented too.